A plane wave traveling along the x-axis in a polystyrene-filled region with πœ€π‘Ÿ = 2.54 has an electric field given by 𝐸𝑦=𝐸0cos(πœ”π‘‘βˆ’π‘˜π‘₯). The frequency is 2.4 GHz, and 𝐸0= 5.0 V/m.

Find the following:

1.1.1. The amplitude and direction of the magnetic field;

1.1.2. The phase velocity;

1.1.3. The wavelength; and

1.1.4. The phase shift between the positions π‘₯1=0.1π‘š and π‘₯2=0.15π‘š


1
Expert's answer
2022-03-02T14:39:07-0500

"H=\sqrt{\frac{\epsilon}{\mu_0}}E=7.1~\frac{kA}m,"

"v=\frac{\omega}k=1.8\cdot 10^8~\frac ms,"

"\lambda=\frac c {\nu}=0.125~m,"

"p=\int_{x_1}^{x_2}Edx=0.23~m."